### v1

@@ Abstract
COVID-19 pandemic has brought to the fore epidemiological models which, though describing a rich variety of behaviors, have previously received little attention in the signal processing literature. During the pandemic, several works successfully leveraged state-of-the-art signal processing strategies to robustly infer epidemiological indicators despite the low quality of COVID-19 data. In the present work, a novel nonstationary autoregressive model is introduced, encompassing, but not reducing to, one of the most popular models for the propagation of viral epidemics. Using a variational framework, penalized likelihood estimators of the parameters of this new model are designed. In practice, the main bottleneck is that the estimation accuracy strongly depends on hyperparameters tuning. Without available ground truth, hyperparameters are selected by minimizing specifically designed data-driven oracles, used as proxy for the estimation error. Focusing on the nonstationary autoregressive Poisson model, the Stein’s Unbiased Risk Estimate formalism is generalized to construct asymptotically unbiased risk estimators based on the derivation of an original autoregressive counterpart of Stein’s lemma. The accuracy of these oracles and of the resulting estimates are assessed through intensive Monte Carlo simulations on synthetic data. Then, elaborating on recent epidemiological models, a novel weekly scaled Poisson model is proposed, enabling to better account for intrinsic variability of the contamination while being robust to reporting errors. Finally, the overall data-driven procedure is particularized to the estimation of COVID-19 reproduction number and exemplified on real COVID-19 infection counts in different countries and at different stages of the pandemic, demonstrating its ability to yield consistent estimates.
@@ 1 Introduction
Most inverse problems of the form (1) are ill-posed, e.g., when $S<T$ leading to non-injective A, hence compromising uniqueness of the quantity of interest given observations under Model (1), or ill-conditioned, e.g., when A is a linear operator with a large conditioning number, inducing numerical instability which, in the presence of noise, translates into prohibitively large estimation variance. Numerous strategies have been designed to address these challenging limitations, which can be divided into two categories: supervised machine learning [10, 64, 37] and unsupervised variational and Bayesian techniques [27, 13, 24, 62]. The present work focuses on the later category, which is particularly adapted when tackling fundamental research problems [14, 50] or newly emerging phenomena [49], for which annotated data are not available. While the negative log-likelihood associated to Model (1) provides a measure of the fidelity of the observations to the model, a regularization term is designed to lift up the ambiguity in the solution of ill-posed and ill-conditioned problems, e.g., leveraging negative log-prior on the quantity of interest [56, 30]. Balancing the fidelity to the data and the regularity constraints then amounts to compute the maximum a posteriori estimate of the quantity of interest as
@@
Related works. Ideally, the hyperparameters would be selected by minimizing the estimation error, that is by choosing
@@
where $\Lambda\subset\mathbb{R}^{L}$ denotes the set of admissible hyperparameters. For linear parametric estimates under additive Gaussian noise hypothesis, the Generalized Cross Validation strategy [29] consists in minimizing a data-dependent criterion constructed as the ratio between the residual sum of squares and the trace of a model-dependent linear operator. Tough, it has been shown that the Generalized Cross Validation function might not have a unique properly defined minimizer [61], leading the Generalized Cross Validation method to fail catastrophically by producing grossly underestimated regularization parameters. Moreover, it is not flexible enough to handle sparsity-inducing penalizations, which are very popular in the inverse problem literature [57]. Among the methods relying on the design of a tractable oracle, Stein’s Unbiased Risk Estimate based strategies, elaborating on the seminal work [60] to construct an approximation of the estimation risk, were initially formulated for i.i.d. Gaussian noise model, but have then been extended to more general noise models [20], notably including a data-dependent Poisson contribution [42, 38, 39]. In the past decades, Stein-based strategies have demonstrated their ability to provide accurate hyperparameter selection in numerous applications, reaching state-of-the-art performance in inverse problem resolution, both in the variational framework as multispectral image deconvolution [2], denoising of force signal in nonlinear physics [50], and more recently in unsupervised deep learning for image denoising [12].
@@
Section 2 first proposes a formal description of a new class of models, namely the generalized nonstationary autoregressive models, providing a general framework encompassing Model (5); then the variational framework is leveraged to design estimators of generalized nonstationary autoregressive models unknown parameters under several commonly encountered noise distributions. The proposed original autoregressive Stein paradigm is developed in Section 3; a novel Stein’s-type lemma is first derived in Section 3.2 for generalized nonstationary autoregressive models involving Poisson noise, and then used to derive a prediction and an estimation unbiased risk estimators in Section 3.3; finally Finite Differences and Monte Carlo strategies are implemented to yield tractable Autoregressive Poisson Unbiased Risk Estimates. The accuracy of the derived risk estimates is supported by intensive numerical simulations on synthetic data, presented in Section 4. Then, in Section 5, a novel weekly scaled Poisson epidemiological model, accounting more precisely for the intrinsic variability of the pathogen propagation while being robust to administrative noise, is introduced. Finally, the Autoregressive Poisson Unbiased Risk Estimate is particularized to this new model to design an original data-driven COVID-19 reproduction number estimator, which is exemplified on real data from different countries worldwide and at different pandemic stages.
@@
2 Nonstationary autoregressive models
@@ 2.1 Observation model
The proposed nonstationary autoregressive model encompasses both the standard autoregressive model of finite order and the epidemiological model introduced in [15] and generalizes them in several directions. First, the memory term $\Psi_{t}$ is not necessarily a linear function, it is only assumed to be causal, that is depending only on past observations, and smooth. Second, observations are no longer restricted to follow an independent Gaussian distribution with constant variance, but instead any distribution with prescribed mean, possibly depending on additional time-varying parameters. Finally, and this is the major originality, observations are externally driven by a time-varying reproduction coefficient.
Definition 1 (Driven autoregressive model).
Let $T\in\mathbb{N}^{*}$ be a time horizon, $\overline{\boldsymbol{\mathsf{X}}}=(\overline{\mathsf{X}}_{1},\ldots,\overline{\mathsf{X}}_{T})\in\mathbb{R}_{+}^{T}$ a time-varying reproduction coefficient, $\mathsf{Y}_{0}\in\mathbb{R}_{+}^{*}$ an initial state, and for each $t\in\{1,\ldots,T\}$, $\Psi_{t}:\mathbb{R}^{t-1}\rightarrow\mathbb{R}$ a smooth function, with by convention $\Psi_{1}=\mathsf{Y}_{0}$. Observations $\boldsymbol{\mathsf{Y}}=(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{T})$ follow a driven autoregressive model with reproduction coefficient $\overline{\boldsymbol{\mathsf{X}}}$ and memory functions $\{\Psi_{t},\,t=1,\ldots,T\}$ if and only if
@@
Example 1 (Driven linear autoregressive model).
The class of linear driven autoregressive models corresponds to linear functions $\{\Psi_{t},\,t=1,\ldots,T\}$ defined as
@@
Linear driven autoregressive models are widely used in epidemiology, where the sequence $\{\psi_{s}\}_{s=1}^{\tau}$ corresponds to the serial interval distribution, accounting for the randomness of the time delay between primary and secondary infections, and $\overline{\mathsf{X}}_{t}$ corresponds to the effective reproduction number [15, 1]. Figure 1 provides a synthetic example of linear driven autoregressive observations under the Poisson model (11), with $\psi$ corresponding to a discretized Gamma distribution of mean $6.6$ and standard deviation $3.5$ truncated at $\tau=25$, mimicking the serial interval function of COVID-19 [8, 55] used in [15, 1, 49]. The three time periods corresponding to $\overline{\mathsf{X}}_{t}>1$ are represented as light blue areas (first row), and results in three temporally separated bumps in $\mathsf{Y}_{t}$ (second row).
@@
The driven autoregressive model of Definition 1 encompasses the standard autoregressive model of order $\tau\in\mathbb{N}^{*}$ defined as
@@
where $\left(\Xi_{t}\right)_{t\in\mathbb{N}^{*}}$ is a sequence of i.i.d. Gaussian variables of zero mean and variance $\alpha^{2}$, for some $\alpha>0$ [59, Section 2.2]. This standard autogressive model indeed corresponds to the linear model described in Example 1 with constant reproduction coefficient $\overline{\mathsf{X}}_{t}=1$, and Gaussian noise with constant variance $\alpha_{t}^{2}=\alpha^{2}$.
It is worth noting that the present work focuses on autoregressive processes which are driven by an external unknown and time-varying reproduction coefficient, which constitutes a paradigm drastically different from the thoroughly studied autoregressive processes, preventing from using the standard tool described, e.g., in [59].
@@
Numerous probability distributions $\mathcal{B}$ are encountered in the inverse problem literature [22, 19, 35, 38, 43]. Three major representative examples adapted to the driven autoregressive model introduced in Definition 1 are: i) the additive Gaussian noise of variance $\alpha_{t}^{2}$
@@
where $\mathcal{G}\left(\alpha,\mathsf{U}\right)$ refers to the Gamma distribution of shape parameter $\alpha$ and scale parameter $\mathsf{U}$ [^3] By convention, whatever $\alpha>0$, if $\mathsf{U}\leq 0$, $\mathcal{P}(\mathsf{U})$ and $\mathcal{G}\left(\alpha,\mathsf{U}\right)$ are Dirac distributions, i.e., $\mathsf{Y}_{t}=0$ deterministically.
Figure 1: Driven Poisson autoregressive data. The process $\mathsf{Y}_{t}$ (top plot, solid curve) follows Model (7) with piecewise linear reproduction coefficient $\mathsf{X}_{t}$ (bottom row, deep blue curve); linear memory functions (8) with horizon $\tau=25$; Poisson degradation (11), with constant scale parameter $\alpha_{t}\equiv 10^{3}$; initialized at $\mathsf{Y}_{0}=10^{4}$ and run for $T=300$ steps. Blue areas corresponds to $\mathsf{X}_{t}\geq 1$, driving exponential growth of $\mathsf{Y}_{t}$.
@@ 2.2 Variational estimators
$-\ln\left(\mathbb{P}(\boldsymbol{\mathsf{Y}}\lvert\boldsymbol{\mathsf{X}};\boldsymbol{\alpha})\right)=\sum_{t=1}^{T}\mathsf{d}_{\mathsf{KL}}\left(\frac{\mathsf{Y}_{t}}{\alpha_{t}}\left\lvert\frac{\mathsf{X}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}{\alpha_{t}}\right.\right),$ (15)
where $\mathsf{d}_{\mathsf{KL}}(\mathsf{Y}\lvert\mathsf{U})=$
$\left\{\begin{array}[]{ll}\mathsf{Y}\ln\left(\frac{\mathsf{Y}}{\mathsf{U}}\right)+\mathsf{U}-\mathsf{Y}&\text{if }\,\mathsf{Y}>0,\,\mathsf{U}>0\\ \mathsf{U}&\text{if }\,\mathsf{Y}=0,\,\mathsf{U}\geq 0\\ \infty&\text{otherwise.}\end{array}\right.$ (19)
Finally under the multiplicative Gamma noise Model (12), the discrepancy is, up to a term independent of $\mathsf{U}$, the Itakura-Saito divergence [22]
$-\ln\left(\mathbb{P}(\boldsymbol{\mathsf{Y}}\lvert\boldsymbol{\mathsf{X}};\boldsymbol{\alpha})\right)=\sum_{t=1}^{T}\mathsf{d}_{\mathsf{IS}}\left(\mathsf{Y}_{t}\lvert\alpha_{t},\frac{\mathsf{X}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}{\alpha_{t}}\right)$ (20)
where $\mathsf{d}_{\mathsf{IS}}(\mathsf{Y}\lvert\alpha,\mathsf{U})=$
$\left\{\begin{array}[]{ll}\frac{\mathsf{Y}}{\mathsf{U}}-\alpha\ln\left(\frac{\mathsf{Y}}{\mathsf{U}}\right)+\ln\left(\Gamma(\alpha)\right)+\ln\left(\mathsf{Y}\right)&\text{if }\,\mathsf{Y}>0,\,\mathsf{U}>0\\ \infty&\text{otherwise,}\end{array}\right.$ (23)
@@
For all three models, Gaussian (10), Poisson (11) and Gamma (12), provided that $\Psi_{t}(\boldsymbol{\mathsf{Y}})$ is positive for all $t\in\{1,\ldots,T\}$, the maximum likelihood estimator (13) writes
@@
An example of Maximum Likelihood estimate on synthetic observations following the Poisson noise model is provided in Figure 1. Due to the presence of noise in the observations (top plot, black solid curve,) the straightforward Maximum Likelihood estimate of the instantaneous reproduction coefficient (bottom plot, light blue curve) suffers from erratic fluctuations compared to ground truth (bottom plot, deep blue curve). Such noisy estimate severely impairs the diagnostic of exponential growth based on $\mathsf{X}_{t}>1$; e.g., around $t=50$, although ground truth is clearly above one, some values $\widehat{\mathsf{X}}^{\mathrm{ML}}_{t}<1$ are observed.
@@
where for each $\ell\in\{1,\ldots,L\}$, $\textbf{L}_{\ell}$ is a linear operator, $q_{\ell}$ a positive exponent, and $\lambda_{\ell}\geq 0$ is a regularization parameter balancing the importance of the $\ell$th constraint with respect to the other constraints in (26) and to the data-fidelity term of Equation (25). Each term of the functional enforces a specific constraint, hence enabling to take into account several regularity and sparsity properties simultaneously. For example, when choosing the discrete Laplacian $\textbf{L}=\textbf{D}_{2}$, $q_{\ell}=1$, favors sparsity of the second order temporal derivative, and hence results in piecewise linear estimates, while $q=2$ yields smooth estimates. The penalized likelihood strategy sketched in Equation (25) is highly flexible and adapts to a large collection of noise models and constraints, hence, by favoring a priori behavior, it has the ability to provide consistent and accurate regularized estimates. The excellent performances of variational estimators are at the price of a cautious fine-tuning of the regularization parameters associated to each term of the penalization. Not only this task is very tough to perform manually but also, and more importantly, in practice ground truth is not available and it is necessary to resort to data-driven oracles to approach optimal hyperparameters.

### v3

@@ Abstract
COVID-19 pandemic has brought to the fore epidemiological models which, though describing a wealth of behaviors, have previously received little attention in the signal processing literature. In this work, a generalized time-varying autoregressive model is considered, encompassing, but not reducing to, a state-of-the-art model of viral epidemics propagation. The time-varying parameter of this model is estimated via the minimization of a penalized likelihood estimator. A major challenge is that the estimation accuracy strongly depends on hyperparameters fine-tuning. Without available ground truth, hyperparameters are selected by minimizing specifically designed data-driven oracles, used as proxy for the estimation error. Focusing on the time-varying autoregressive Poisson model, the Stein’s Unbiased Risk Estimate formalism is generalized to construct asymptotically unbiased risk estimators based on the derivation of an original autoregressive counterpart of Stein’s lemma. The accuracy of these oracles and of the resulting estimates are assessed through intensive Monte Carlo simulations on synthetic data. Then, elaborating on recent epidemiological models, a novel weekly scaled Poisson model is proposed, better accounting for intrinsic variability of the contamination while being robust to reporting errors. Finally, the overall data-driven procedure is particularized to the estimation of COVID-19 reproduction number demonstrating its ability to yield very consistent estimates.
@@ 1 Introduction
Most inverse problems of the form (1) are ill-posed, e.g., when $S<T$ leading to non-injective A, hence compromising uniqueness of the quantity of interest given observations under Model (1), or ill-conditioned, e.g., when A is a linear operator with a large conditioning number, inducing numerical instability which, in the presence of noise, translates into prohibitively large estimation variance. Numerous strategies have been designed to address these challenging limitations, including variational approaches computing a deterministic solution [75, 80, 24, 15, 28, 76], Bayesian techniques well-suited to probability distribution probing [72, 32, 52, 29], and the recently developed supervised machine learning strategies relying on abundant annotated data [11, 79, 46]. The present work focuses on unsupervised variational approaches, which are particularly adapted when tackling fundamental research problems [16, 60] or newly emerging phenomena [59], for which complex physical models have been proposed but with no annotated data available. While the negative log-likelihood associated to Model (1) provides a measure of the fidelity of the observations to the model, a regularization term is designed to lift the ambiguity in the solution of ill-posed and ill-conditioned problems, e.g., leveraging negative log-prior on the quantity of interest [68, 35]. Balancing the fidelity to the data and the regularity constraints then amounts to compute the maximum a posteriori estimate of the quantity of interest as
@@
Related works. Ideally, the hyperparameters would be selected by minimizing the estimation error choosing
@@
where $\Lambda\subset\mathbb{R}^{L}$ denotes the set of admissible hyperparameters. For linear parametric estimates under additive Gaussian noise hypothesis, the Generalized Cross Validation strategy [34] consists in minimizing a data-dependent criterion constructed as the ratio between the residual sum of squares and the trace of a model-dependent linear operator. Several extensions have been proposed to extend it to sparsity-inducing estimators [74, 42]. Though, it has been shown that the Generalized Cross Validation function might not have a unique properly defined minimizer [73], leading the Generalized Cross Validation method to fail catastrophically by producing grossly underestimated regularization parameters. Furthermore and even more importantly, Generalized Cross Validation requires that the degradation $\mathcal{B}$ in (1) is independent and stationary across observations [34] which is a very restrictive hypothesis excluding, e.g., heteroscedastic [39] and correlated noise [61]. The recently proposed Approximate Leave-One-Out (ALO) Cross Validation procedure, while particularly efficient and scaling to high-dimensional problems, also relies on independence and stationarity assumptions [3, 56]. Among the methods relying on the design of a tractable oracle, Stein’s Unbiased Risk Estimate based strategies, elaborating on the seminal work [71] to construct an approximation of the estimation risk, were initially formulated for i.i.d. Gaussian noise model, but have then been extended to more general noise models [23], notably including a data-dependent Poisson contribution [51, 47, 48]. In the past decades, Stein-based strategies have demonstrated their ability to provide accurate hyperparameter selection in numerous applications, reaching state-of-the-art performance in inverse problem resolution, both in the variational framework, e.g., for multispectral image deconvolution [2] or denoising of force signal in nonlinear physics [60], and more recently in unsupervised deep learning for image denoising [14].
@@
Section 2 first proposes a formal description of the generalized time-varying autoregressive models considered, providing a general framework encompassing Model (5); then the variational framework is leveraged to design estimators of generalized time-varying autoregressive models unknown parameters under several commonly encountered noise distributions. The proposed original autoregressive Stein paradigm is developed in Section 3; a novel Stein’s-type lemma is first derived in Section 3.2 for generalized time-varying autoregressive models involving Poisson noise, and then used to derive a prediction and an estimation unbiased risk estimators in Section 3.3; finally Finite Differences and Monte Carlo strategies are implemented to yield tractable Autoregressive Poisson Unbiased Risk Estimates. The accuracy of the derived risk estimates is supported by intensive numerical simulations on synthetic data, presented in Section 4. Then, in Section 5, a novel weekly scaled Poisson epidemiological model, accounting more precisely for the intrinsic variability of the pathogen propagation while being robust to administrative noise, is introduced. Finally, the Autoregressive Poisson Unbiased Risk Estimate is particularized to this new model to design an original data-driven COVID-19 reproduction number estimator, which is exemplified on real data from different countries worldwide and at different pandemic stages.
@@
2 Generalized time-varying autoregressive models
@@ 2.1 Observation model
The present work considers a generalized time-varying autoregressive model encompassing both the standard autoregressive model of finite order and the epidemiological model introduced in [17] while generalizing both of them in several directions. The first major originality is that the model is externally controlled by an unknown deterministic time-varying reproduction coefficient containing the information of interest. This setup drastically contrasts both with the standard time series framework, mostly interest in predicting the behavior of the observations rather than inferring the coefficients [18], and with random coefficient autoregressive models in which the time-varying coefficients are random variable [38, 63]. Second, observations are no longer restricted to follow an independent Gaussian distribution with constant variance, but instead any distribution with prescribed mean, possibly depending on additional time-varying parameters. Importantly the noise is not necessary additive but might be data-dependent or multiplicative. And finally, the memory term $\Psi_{t}$ is not necessarily a linear function, it is only assumed to be causal, that is depending only on past observations, and smooth.
Definition 1 (Generalized time-varying autoregressive model).
Let $T\in\mathbb{N}^{*}$ be a time horizon, $\overline{\boldsymbol{\mathsf{X}}}=(\overline{\mathsf{X}}_{1},\ldots,\overline{\mathsf{X}}_{T})\in\mathbb{R}_{+}^{T}$ a time-varying reproduction coefficient, $\mathsf{Y}_{0}\in\mathbb{R}_{+}^{*}$ an initial state, and for each $t\in\{1,\ldots,T\}$, $\Psi_{t}:\mathbb{R}^{t-1}\rightarrow\mathbb{R}$ a smooth function, with by convention $\Psi_{1}=\mathsf{Y}_{0}$. Observations $\boldsymbol{\mathsf{Y}}=(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{T})$ follow a generalized time-varying autoregressive model with reproduction coefficient $\overline{\boldsymbol{\mathsf{X}}}$ and memory functions $\{\Psi_{t},\,t=1,\ldots,T\}$ if and only if
@@
Example 1 (Linear time-varying autoregressive model).
The class of linear time-varying autoregressive models corresponds to linear functions $\{\Psi_{t},\,t=1,\ldots,T\}$ defined as
@@
Linear time-varying autoregressive models are widely used in epidemiology, where the sequence $\{\psi_{s}\}_{s=1}^{\tau}$ corresponds to the serial interval distribution, accounting for the randomness of the time delay between primary and secondary infections, and $\overline{\mathsf{X}}_{t}$ corresponds to the effective reproduction number [17, 1]. Figure 1(b) provides a synthetic example of linear time-varying autoregressive observations under the Poisson model (11), with $\psi$ corresponding to a discretized Gamma distribution of mean $6.6$ and standard deviation $3.5$ truncated at $\tau=25$, mimicking the serial interval function of COVID-19 [9, 66] used in [17, 1, 59]. The three time periods corresponding to $\overline{\mathsf{X}}_{t}>1$ are represented as light blue areas (first row), and results in three temporally separated bumps in $\mathsf{Y}_{t}$ (second row).
@@
The generalized time-varying autoregressive model of Definition 1 encompasses the standard autoregressive model of order $\tau\in\mathbb{N}^{*}$ defined as
@@
where $\left(\Xi_{t}\right)_{t\in\mathbb{N}^{*}}$ is a sequence of i.i.d. Gaussian variables of zero mean and variance $\alpha^{2}$, for some $\alpha>0$ [70, Section 2.2]. This standard autogressive model indeed corresponds to the linear model described in Example 1 with constant reproduction coefficient $\overline{\mathsf{X}}_{t}=1$, and Gaussian noise with constant variance $\alpha_{t}^{2}=\alpha^{2}$.
It is worth noting that the present work focuses on time-varying autoregressive processes which are driven by an external unknown and time-varying reproduction coefficient, which constitutes a paradigm drastically different from the thoroughly studied autoregressive processes, preventing from using the standard tool described, e.g., in [70, 18].
@@
Numerous probability distributions $\mathcal{B}$ are encountered in the inverse problem literature [26, 22, 43, 47, 52]. Three major representative examples adapted to the generalized time-varying autoregressive model introduced in Definition 1 are: i) the additive Gaussian noise of variance $\alpha_{t}^{2}$
@@
where $\mathcal{G}\left(\alpha,\mathsf{U}\right)$ refers to the Gamma distribution of shape parameter $\alpha$ and scale parameter $\mathsf{U}$ [^3] By convention, whatever $\alpha>0$, if $\mathsf{U}\leq 0$, $\mathcal{P}(\mathsf{U})$ and $\mathcal{G}\left(\alpha,\mathsf{U}\right)$ are Dirac distributions, i.e., $\mathsf{Y}_{t}=0$ deterministically. Both the Gaussian and Gamma distributions are well-suited to described continuous quantities. The seminal additive Gaussian noise model is much used in biomedical signal analysis, for example to analyze electroencephalogram (EEG) recordings [84, 12], and in communications [83, 77], while the multiplicative Gamma noise model has been used in financial econometrics [45]. Data-dependent Poisson noise is mostly used for count time series in which the observations can take only discrete, typically integer, values [63], with applications in epidemiology [81, 17] and in climate studies, for example to analyze changepoints in the North Atlantic tropical cyclone record [67]. Figure 1 shows on the top row generalized time-varying autoregressive data under Model (7) with respectively additive Gaussian noise 1(a), data-dependent Poisson noise 1(b), and multiplicative Gamma noise 1(c) with prescribed time-varying reproduction coefficient, displayed in dark blue on the bottom row. For all three examples, the function $\psi$ shaping the memory term corresponds to a discretized Gamma distribution of mean $6.6$ and standard deviation $3.5$ truncated at $\tau=25$, mimicking the serial interval function of COVID-19 [9, 66].
@@ 2.2 Variational estimators
$-\ln\left(\mathbb{P}(\boldsymbol{\mathsf{Y}}\lvert\boldsymbol{\mathsf{X}};\boldsymbol{\alpha})\right)=\sum_{t=1}^{T}\mathsf{d}_{\mathsf{KL}}\left(\frac{\mathsf{Y}_{t}}{\alpha_{t}}\left\lvert\frac{\mathsf{X}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}{\alpha_{t}}\right.\right),\quad\mathsf{d}_{\mathsf{KL}}(\mathsf{Y}\lvert\mathsf{U})=\left\{\begin{array}[]{ll}\mathsf{Y}\ln\left(\frac{\mathsf{Y}}{\mathsf{U}}\right)+\mathsf{U}-\mathsf{Y}&\text{if }\,\mathsf{Y}>0,\,\mathsf{U}>0\\ \mathsf{U}&\text{if }\,\mathsf{Y}=0,\,\mathsf{U}\geq 0\\ \infty&\text{otherwise.}\end{array}\right.$ (18)
Finally, under the multiplicative Gamma noise Model (12), the discrepancy is, up to a term independent of $\mathsf{U}$, the Itakura-Saito divergence [26]
$-\ln\left(\mathbb{P}(\boldsymbol{\mathsf{Y}}\lvert\boldsymbol{\mathsf{X}};\boldsymbol{\alpha})\right)=\sum_{t=1}^{T}\mathsf{d}_{\mathsf{IS}}\left(\mathsf{Y}_{t}\lvert\alpha_{t},\frac{\mathsf{X}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}{\alpha_{t}}\right),\quad\mathsf{d}_{\mathsf{IS}}(\mathsf{Y}\lvert\alpha,\mathsf{U})=\left\{\begin{array}[]{ll}\frac{\mathsf{Y}}{\mathsf{U}}-\alpha\ln\left(\frac{\mathsf{Y}}{\mathsf{U}}\right)+\ln\left(\Gamma(\alpha)\mathsf{Y}\right)&\text{if }\,\mathsf{Y}>0,\,\mathsf{U}>0\\ \infty&\text{otherwise,}\end{array}\right.$ (21)
@@
For all three models, Gaussian (10), Poisson (11) and Gamma (12), provided that $\Psi_{t}(\boldsymbol{\mathsf{Y}})$ is nonzero for all $t\in\{1,\ldots,T\}$, the maximum likelihood estimator (13) writes
@@
The Maximum Likelihood estimates (13) of the underlying reproduction number under all three models are displayed in light blue on the bottom row of Figure 1. Due to the presence of noise in the observations (top plot, black solid curve) the straightforward Maximum Likelihood estimate of the instantaneous reproduction coefficient (bottom plot, light blue curve) suffers from erratic fluctuations compared to ground truth (bottom plot, deep blue curve). Such noisy estimate severely impairs the diagnostic of exponential growth based on $\mathsf{X}_{t}>1$; e.g., in Figure 1(b), bottom plot, around $t=50$, although ground truth is clearly above one, some values $\widehat{\mathsf{X}}^{\mathrm{ML}}_{t}<1$ are observed.
@@
where for each $\ell\in\{1,\ldots,L\}$, $\textbf{L}_{\ell}$ is a linear operator, $q_{\ell}$ a positive exponent, and $\lambda_{\ell}\geq 0$ is a regularization parameter balancing the importance of the $\ell$th constraint with respect to the other constraints in (24) and to the data-fidelity term of Equation (23). Each term of the functional enforces a specific constraint, hence enabling to take into account several regularity and sparsity properties simultaneously. For example, when choosing the discrete Laplacian $\textbf{L}=\textbf{D}_{2}$, $q_{\ell}=1$, favors sparsity of the second order temporal derivative, and hence results in piecewise linear estimates, while $q=2$ yields smooth estimates. The penalized likelihood strategy sketched in Equation (23) is highly flexible and adapts to a large collection of noise models and constraints, hence, by favoring a priori behavior, it has the ability to provide consistent and accurate regularized estimates. The excellent performance of variational estimators comes at the price of a cautious fine-tuning of the regularization parameters associated to each term of the penalization. Not only is this task very tough to perform manually but also, and more importantly, in practice ground truth is not available, and it is necessary to resort to data-driven oracles to approach optimal hyperparameters.
(a) Additive Gaussian Model (10)
(b) Data-dependent Poisson Model (11)
(c) Multiplicative Gamma Model (12)
Figure 1: Generalized time-varying autoregressive data and Maximum Likelihood reproduction coefficient estimates. The processes $\mathsf{Y}_{t}$, displayed as solid black curves on top row, follow Model (7) with piecewise linear reproduction coefficient $\mathsf{X}_{t}$, displayed as dark blue curve on the bottom row; linear memory functions (8) with horizon $\tau=25$; initialized at $\mathsf{Y}_{0}=10^{4}$ and run for $T=300$ steps. The associated memory terms $\Psi_{t}(\boldsymbol{\mathsf{Y}})$ are displayed as dashed black curves on the top row. Blue areas on the bottom row corresponds to $\mathsf{X}_{t}\geq 1$ inducing exponential growth of $\mathsf{Y}_{t}$. The Maximum likelihood estimates (13) of the underlying reproduction coefficient under each model are displayed as light blue curves on the bottom row. (a) Additive Gaussian noise (10) with constant standard deviation $\alpha_{t}\equiv 10^{5}$. (b) Poisson degradation (11), with constant scale parameter $\alpha_{t}\equiv 10^{3}$. (c) Gamma degradation (12), with constant shape parameter $\alpha_{t}\equiv 20$.